# Causal machine learning methods and use of sample splitting in settings with high-dimensional confounding

### v1

@@ Abstract
Observational epidemiological studies commonly seek to estimate the causal effect of an exposure on an outcome. Adjustment for potential confounding bias in modern studies is challenging due to the presence of high-dimensional confounding, induced when there are many confounders relative to sample size, or complex relationships between continuous confounders and exposure and outcome. As a promising avenue to overcome this challenge, doubly robust methods (Augmented Inverse Probability Weighting (AIPW) and Targeted Maximum Likelihood Estimation (TMLE)) enable the use of data-adaptive approaches to fit the two models they involve. A challenge when applying these methods is that biased standard errors may result when the data-adaptive approaches used are very complex. The coupling of doubly robust methods with cross-fitting has been proposed to tackle this. Despite advances, limited evaluation, comparison, and guidance are available on the implementation of AIPW and TMLE with data-adaptive approaches and cross-fitting in realistic settings where high-dimensional confounding is present. Motivated by an early-life cohort study, we conducted an extensive simulation study to compare the relative performance of AIPW and TMLE using data-adaptive approaches in estimating the average causal effect (ACE) and evaluated the benefits of using cross-fitting with a varying number of folds, as well as the impact of using a reduced versus full (larger, more diverse) library in the Super Learner ensemble learning approach used for the data-adaptive models. A range of scenarios in terms of data generation, and sample size were considered. We found that AIPW and TMLE performed similarly in most cases for estimating the ACE, but TMLE was more stable. Cross-fitting improved the performance of both methods, with the number of folds a less important consideration. Using a full Super Learner library was important to reduce bias and variance in the complex scenarios typical of modern health research studies. We illustrated the methods and practical value of our findings in a case study that investigated the ACE of early life inflammation on later cardiovascular risk.
Keywords Doubly robust  $\cdot$ High-dimensional confounding  $\cdot$ Augmented Inverse Probability Weighting  $\cdot$ Targeted Maximum Likelihood Estimation  $\cdot$ Machine learning  $\cdot$ Sample splitting  $\cdot$ Causal inference  $\cdot$ Cross-fitting
@@ 1 Introduction
Commonly used methods for the estimation of causal effects, such as g-computation or inverse probability weighting (IPW) are singly robust (SR), meaning they rely upon a single model being correctly specified (Naimi et al., 2023; Schuler and Rose, 2017). G-computation involves the fitting of an outcome model (conditional expectation of the outcome given the exposure and confounders), and IPW an exposure model (probability of exposure given confounders). In the context of high-dimensional confounding, such methods are at risk of misspecification bias because the single model is commonly based on simplistic parametric functions (Naimi et al., 2023; Zivich and Breskin, 2021). Specifically, in settings with a large number of confounders (or higher-order terms) and limited sample size, it can be challenging to apply the methods without some form of variable selection or model simplification, and attempting to do this naively can impact the validity of inferences (Dukes and Vansteelandt, 2020; Leeb and Pöscher, 2006). Indeed, many studies reduce to a restricted, manageable adjustment set that may not be sufficient as an adjustment set resulting in bias. Additionally, the failure to capture complex nonlinear relationships between variables can potentially lead to further bias (Díaz, 2019; Schuler and Rose, 2017; VanderWeele, 2019; Groenwold et al., 2009).
@@
To overcome these concerns, it is tempting to consider incorporating data-adaptive approaches to flexibly fit the single model required for SR methods. We define data-adaptive approaches as those where the modelling approach or algorithm used to fit the model is capable of learning from the data, in the sense that the full functional form is not fixed but can adapt. Data-adaptive approaches include machine learning (ML) approaches, and the terms are frequently used interchangeably (Rose and van der Laan, 2011, pg.44), although data-adaptive can also refer to parametric variable selection approaches like stepwise regression, an approach that we do not consider here. Unfortunately, the use of data-adaptive approaches with SR methods has been shown to be problematic, leading to bias in point estimates and with no valid approach to obtain standard errors (Van der Laan and Rubin, 2006; Chernozhukov et al., 2017; Naimi et al., 2023; Balzer and Westling, 2023). For example, for g-computation with data-adaptive approaches, the non-parametric bootstrap has been shown to be invalid for variance estimation (Bickel et al., 1997). These issues arise due to slow convergence rates which lead to bias and challenges in the construction of CIs (van der Vaart, 2014), and underestimated variance and under-coverage have been reported, signalling that misleading inference is a concern (Naimi et al., 2023; Balzer and Westling, 2023).
@@
As promising alternatives to SR methods, Augmented Inverse Probability Weighting (AIPW) (Bang and Robins, 2005; Robins et al., 1994; Glynn and Quinn, 2010) and Targeted Maximum Likelihood Estimation (TMLE) (Rose and van der Laan, 2011; Lendle et al., 2013; Rose and Rizopoulos, 2019) involve the fitting of both an outcome and an exposure model. AIPW and TMLE have attractive properties, including that they provide consistent estimation if at least one of the two models is consistently estimated, which is why they are referred to as doubly robust (DR). These methods also achieve optimal non-parametric efficiency if both models are consistently estimated (Luque-Fernandez et al., 2018; Robins et al., 2007). The DR methods are increasingly being used to estimate causal effects because of these properties, providing the opportunity to reduce the degree of bias due to model misspecification in both low and high-dimensional confounding settings. In contrast to SR methods, DR methods can validly incorporate the use of data-adaptive approaches to fit both models under some conditions, potentially further reducing the risk of misspecification bias as well as providing a way to avoid inappropriate a priori variable selection in high-dimensional confounding settings. Data-adaptive estimators typically have non-standard asymptotic behaviour, meaning that they are often non-normal, and their standard errors do not converge to zero at a rate of $1/\sqrt{n}$ (Zivich and Breskin, 2021), meaning that slower convergence of models fitted via data-adaptive means is common. Although this is problematic for incorporating these approaches into SR methods as previously mentioned, it is not the case for incorporating them into DR estimators, which remain asymptotically normal and root-n-consistent under some conditions (Naimi et al., 2023).
@@
A challenge in the application of DR methods with data-adaptive approaches is the estimation of standard errors. Indeed, it has been found that biased standard error estimates may result when using the DR methods with some very complex and flexible data-adaptive approaches (Kennedy, 2016; Chernozhukov et al., 2018; Zheng and van der Laan, 2011; Newey and Robins, 2018), although it should be noted that this still offers benefits over SR methods, where even obtaining standard errors can be problematic (Naimi et al., 2023). Sample splitting, whereby the outcome and exposure models are fit in subsets (folds) of the data and the causal effect estimate is obtained from the remaining data, has been proposed to improve the estimation of standard errors in the DR method context (Chernozhukov et al., 2018; Zheng and van der Laan, 2011).
@@
Despite advances, there is limited evaluation, comparison, and guidance on the implementation of what we will henceforth refer to as “causal machine learning” (CML) (AIPW and TMLE with data-adaptive approaches) with sample splitting in realistic settings encountered in modern observational studies with high-dimensional confounding (Meng and Huang, 2022). Most methodological studies evaluate and compare the methods with and without sample splitting in less realistic settings, with very few or binary confounders only (Naimi et al., 2023; Zivich and Breskin, 2021). Empirical studies applying these methods often have large sample size, but many real-world studies have sample size limitations. In general, empirical studies applying these methods have been limited to settings where high-dimensional confounding has not been considered, when perhaps it should have been (Lendle et al., 2013; Kreif and DiazOrdaz, 2019; Decruyenaere et al., 2020; Herrera et al., 2018). To our knowledge no studies have comprehensively evaluated and compared AIPW and TMLE, with and without sample splitting, whilst considering varying sample sizes in the high-dimensional setting. In addition, we are not aware of any studies that have evaluated how the number of folds used in sample splitting may affect the performance of AIPW and TMLE.
@@
Here, motivated by a real-world case study, we conduct an extensive simulation study to address these gaps in the context of estimating the average causal effect (ACE), to provide practical guidance on the application of CML methods in realistic settings. The manuscript is organised as follows. Section 2 introduces the motivating case study. Section 3 outlines relevant notation and assumptions and Section 4 provides details on AIPW and TMLE for the estimation of causal effects. Section 5 outlines key considerations for the implementation and evaluation of DR methods.
Section 6 and 7 describe the simulation study design and results. In Section 8, to illustrate, we apply the methods to the case study, following which in Section 9 we discuss our findings and their implications.
@@ 2 Overview of the motivating case study
The motivating example draws on data from the Barwon Infant Study (BIS), a birth cohort study with antenatal recruitment, conducted in the south-east of Australia (Vuillermin et al., 2015). Infants included in the study were reviewed at multiple time points in early life (including at birth, 12 months, and 4 years of age). Further details regarding eligibility and recruitment criteria for BIS are detailed elsewhere (Vuillermin et al., 2015).
@@
One focus area for research in BIS is the developmental origins of cardiovascular disease (CVD). There is growing evidence for a role of inflammation in CVD (Sorriento and Iaccarino, 2019; DeWeerdt, 2021). Chiesa et al. (2022) found that inflammation (as measured by a biomarker called Glycoprotein acetyls (GlycA) obtained directly from blood samples) was associated with adverse cardiovascular profiles in adolescence, and also predicted future risk. In adults, pulse wave velocity (PWV), a measure of arterial stiffness (higher PWV can indicate greater arterial stiffness), has been shown to predict CVD events (Sutton-Tyrrell et al., 2005; Sequí-Domínguez et al., 2020). It is therefore of interest to examine the effect of early life inflammation on PWV at a later age in childhood (post infancy). The motivating example focusses specifically on the following research question: What is the effect of early life inflammation, as measured by GlycA in 1-year-old infants, on PWV at 4 years of age? We standardise the outcome (PWV) in the BIS data and consider GlycA as a dichotomous exposure, indicating either high or low inflammation. The specific GlycA concentration threshold above which inflammation is considered high in infants remains unclear in the literature. We somewhat arbitrarily dichotomised GlycA at its 75th percentile to provide a focus on the common setting of a dichotomous exposure for the purpose of our methodological investigation.
@@
To address the proposed causal question, one must consider the potential confounding role of background factors (e.g., demographic, environmental, familial, and perinatal) and other metabolomic factors, as depicted in the directed acyclic graph (DAG) of Figure LABEL:fig:BIS_DAG developed using subject-matter expertise. In BIS, metabolomic measures at 1 year of age and PWV at 4 years of age were only obtained for a subset of the participants by design. Details on missing data are provided in Table S1 of Section 1 of the Supplementary material. Only participants with complete data (no missing data on any of the relevant variables, n=252) were included in the case study, as evaluation of challenges associated with the presence of missing data were not considered here. Characteristics of the infants included are outlined in Table LABEL:table:BIS_descriptive. Here, we have clearly identified a problem of high-dimensional confounding because we have a sample size of 252 and up to 87 potential confounders, with 78 of these being continuous confounders.
@@ 3.1 Causal estimand
We focus on the average causal effect (ACE), defined as $E[Y^{X=1}]-E[Y^{X=0}]$, interpreted as the difference in average potential outcome in the population of interest when (a) everyone is exposed versus when (b) everyone is unexposed (Hernan, 2004; Rubin, 2005). In the motivating example, this is the difference in average PWV that would be seen at 4 years of age if everyone in the target population were set to have high versus low GlycA at 1 year of age.
@@ 4 Doubly robust methods
Under the assumptions outlined in Section 3, the ACE can be estimated using DR methods. We focus on two DR methods, AIPW and TMLE. For both of these methods, models are fitted for the conditional expectation of the outcome $Y$ given the exposure $X$ and confounders $W$, $E[Y|X,W]$ (the outcome model) and for the propensity score, ${P}(X=1|W)$ (the exposure model). The outcome and exposure models are often referred to as nuisance models because they are not of intrinsic interest but instead are used within AIPW and TMLE to estimate the target parameter (Daniel, 2018).
DR estimators are obtained by determining and then utilising an efficient influence function (EIF), which has a unique form that is determined by the target parameter of interest, in our case the ACE (Kennedy, 2016). DR estimators are constructed in a manner that allows them to attempt to correct for the bias (termed plug-in bias) that is induced when using g-computation with the data-adaptive approaches. Further details regarding derivation and interpretation of the EIF, as well as plug-in bias can be found elsewhere (Hines et al., 2022).
@@ 4.1 Augmented Inverse Probability Weighting (AIPW)
AIPW is often referred to as a one-step correction or one-step estimation approach (Benkeser et al., 2017) and is based on directly subtracting an estimate of the bias term from the g-computation estimator. Based on the fitted outcome model, predicted values, $\hat{E}_{1}(w)=\hat{E}[Y|X=1,W=w]$ and $\hat{E}_{0}(w)=\hat{E}[Y|X=0,W=w]$ are obtained for each record, for $X=1$ and $X=0$ respectively. The estimate of the ACE is then calculated as
@@
where $\hat{P}(w)=\hat{P}(X=1|W=w)$ is an estimate of the propensity score based on the fitted exposure model. The estimated variance of the AIPW estimator is obtained as follows (Wager, 2022; Chernozhukov et al., 2018):

### v2

# Susan Ellul

@@ Abstract
[Abstract]Observational epidemiological studies commonly seek to estimate the causal effect of an exposure on an outcome. Adjustment for potential confounding bias in modern studies is challenging due to the presence of high-dimensional confounding, which occurs when there are many confounders relative to sample size or complex relationships between continuous confounders and exposure and outcome. Despite recent advances, limited evaluation, and guidance are available on the implementation of doubly robust methods, Augmented Inverse Probability Weighting (AIPW) and Targeted Maximum Likelihood Estimation (TMLE), with data-adaptive approaches and cross-fitting in realistic settings where high-dimensional confounding is present. Motivated by an early-life cohort study, we conducted an extensive simulation study to compare the relative performance of AIPW and TMLE using data-adaptive approaches in estimating the average causal effect (ACE). We evaluated the benefits of using cross-fitting with a varying number of folds, as well as the impact of using a reduced versus full (larger, more diverse) library in the Super Learner ensemble learning approach used for implementation. We found that AIPW and TMLE performed similarly in most cases for estimating the ACE, but TMLE was more stable. Cross-fitting improved the performance of both methods, but was more important for estimation of standard error and coverage than for point estimates, with the number of folds a less important consideration. Using a full Super Learner library was important to reduce bias and variance in complex scenarios typical of modern health research studies.
[…]
keywords:
Causal inference, Doubly robust, High-dimensional confounding, Augmented Inverse Probability Weighting, Targeted Maximum Likelihood Estimation, Cross-fitting
articletype: Research Articlefootnotetext: Abbreviations: ACE, average causal effect; AIPW, Augmented inverse probability weighting; CML, causal machine learning; TMLE, Targeted maximum likelihood estimation.
@@ 1 Introduction
Commonly used methods for the estimation of causal effects, such as g-computation or inverse probability weighting (IPW) are singly robust (SR), meaning they rely upon a single model being correctly specified. 1, 2 G-computation involves the fitting of an outcome model (conditional expectation of the outcome given the exposure and confounders), and IPW an exposure model (probability of exposure given confounders). In the context of high-dimensional confounding, such methods are at risk of misspecification bias because the single model is commonly based on simplistic parametric functions. 1, 3 Specifically, in settings with a large number of confounders (or higher-order terms) and limited sample size, it can be challenging to apply the methods without some form of variable selection or model simplification, and attempting to do this naively can impact the validity of inferences. 4, 5 Indeed, many studies reduce to a restricted, manageable adjustment set that may not be sufficient as an adjustment set resulting in bias. Additionally, the failure to capture complex nonlinear relationships between variables can potentially lead to further bias.2, 6, 7, 8
@@
To overcome these concerns, it is tempting to consider incorporating data-adaptive approaches to flexibly fit the single model required for SR methods. We define data-adaptive approaches as those where the modelling approach or algorithm used to fit the model is capable of learning from the data, in the sense that the full functional form is not fixed but can adapt. Data-adaptive approaches include machine learning (ML) approaches, and the terms are frequently used interchangeably 9 (p 44), although data-adaptive can also refer to parametric variable selection approaches like stepwise regression, an approach that we do not consider here. Unfortunately, the use of data-adaptive approaches with SR methods has been shown to be problematic, leading to bias in point estimates and with no valid approach to obtain standard errors. 10, 11, 1, 12 For example, for g-computation with data-adaptive approaches, the non-parametric bootstrap has been shown to be invalid for variance estimation. 13 These issues arise due to slow convergence rates which lead to bias and challenges in the construction of CIs 14, and underestimated variance and under-coverage have been reported, signalling that misleading inference is a concern. 1, 12
@@
As promising alternatives to SR methods, Augmented Inverse Probability Weighting (AIPW) 15, 16, 17 and Targeted Maximum Likelihood Estimation (TMLE) 9, 18, 19 involve the fitting of both an outcome and an exposure model. AIPW and TMLE have attractive properties, including that they provide consistent estimation if at least one of the two models is consistently estimated, which is why they are referred to as doubly robust (DR). These methods also achieve optimal non-parametric efficiency if both models are consistently estimated. 20, 21 The DR methods are increasingly being used to estimate causal effects because of these properties, providing the opportunity to reduce the degree of bias due to model misspecification in both low and high-dimensional confounding settings. In contrast to SR methods, DR methods can validly incorporate the use of data-adaptive approaches to fit both models under some conditions, potentially further reducing the risk of misspecification bias as well as providing a way to avoid inappropriate a priori variable selection in high-dimensional confounding settings. Data-adaptive estimators typically have non-standard asymptotic behaviour, meaning that they are often non-normal, and their standard errors do not converge to zero at a rate of $1/\sqrt{n}$. 3 Although this is problematic for incorporating these approaches into SR methods as previously mentioned, it is not the case for incorporating them into DR estimators, which remain asymptotically normal and root-n-consistent under some conditions. 1
@@
A challenge in the application of DR methods with data-adaptive approaches is the estimation of standard errors. Indeed, it has been found that biased standard error estimates may result when using the DR methods with some very complex and flexible data-adaptive approaches 22, 23, 24, 25, although it should be noted that this still offers benefits over SR methods, where even obtaining standard errors can be problematic. 1 Sample splitting, whereby the outcome and exposure models are fit in subsets (folds) of the data and the causal effect estimate is obtained from the remaining data, has been proposed to improve the estimation of standard errors in the DR method context.23, 24
@@
Despite advances, there is limited evaluation, comparison, and guidance on the implementation of what we will henceforth refer to as “causal machine learning” (CML) (AIPW and TMLE with data-adaptive approaches) with sample splitting in realistic settings encountered in modern observational studies with high-dimensional confounding. 26 Most methodological studies evaluate and compare the methods with and without sample splitting in less realistic settings, with very few or binary confounders only. 1, 3 Empirical studies applying these methods often have large sample size, but many real-world studies have sample size limitations. In general, empirical studies applying these methods have been limited to settings where high-dimensional confounding has not been considered, when perhaps it should have been. 18, 27, 28, 29 To our knowledge no studies have comprehensively evaluated and compared AIPW and TMLE, with and without sample splitting, whilst considering varying sample sizes in the high-dimensional setting. In addition, we are not aware of any studies that have evaluated how the number of folds used in sample splitting may affect the performance of AIPW and TMLE.
@@
Here, motivated by a real-world case study, we conduct an extensive simulation study to address these gaps in the context of estimating the average causal effect (ACE), to provide practical guidance on the application of CML methods in realistic settings. The manuscript is organised as follows. Section 2 introduces the motivating case study. Section 3 outlines relevant notation and assumptions and Section 4 provides details on AIPW and TMLE for the estimation of causal effects. Section 5 outlines key considerations for the implementation and evaluation of DR methods. Section 6 and 7 describe the simulation study design and results. In Section 8, to illustrate, we apply the methods to the case study, following which in Section 9 we discuss our findings and their implications.
@@ 2 Overview of the motivating case study
The motivating example draws on data from the Barwon Infant Study (BIS), a birth cohort study with antenatal recruitment, conducted in the south-east of Australia. 30 Infants included in the study were reviewed at multiple time points in early life (including at birth, 12 months, and 4 years of age). Further details regarding eligibility, recruitment criteria and measures obtained for BIS are provided elsewhere. 30
@@
One focus area for research in BIS is the developmental origins of cardiovascular disease (CVD). There is growing evidence for a role of inflammation in CVD. 31, 32 Chiesa et al 33 found that inflammation (as measured by a biomarker called Glycoprotein acetyls (GlycA) obtained directly from blood samples) was associated with adverse cardiovascular profiles in adolescence, and also predicted future risk. In adults, pulse wave velocity (PWV), a measure of arterial stiffness (higher PWV can indicate greater arterial stiffness), has been shown to predict CVD events. 34, 35 It is therefore of interest to examine the effect of early life inflammation on PWV at a later age in childhood (post infancy). The motivating example focusses specifically on the following research question: What is the effect of early life inflammation, as measured by GlycA in 1-year-old infants, on PWV at 4 years of age? We standardise the outcome (PWV) in the BIS data, and consider GlycA as a dichotomous exposure, indicating either high or low inflammation. We dichotomise the exposure to make results relevant for the majority of applications that currently focus on binary exposures. The specific GlycA concentration threshold above which inflammation is considered high in infants remains unclear in the literature. We somewhat arbitrarily dichotomised GlycA at its 75th percentile.
@@
To address the proposed causal question, one must consider the potential confounding role of background factors (e.g., demographic, environmental, familial, and perinatal) and other metabolomic factors, as depicted in the directed acyclic graph (DAG) of Figure 1 developed using subject-matter expertise. In BIS, metabolomic measures at 1 year of age and PWV at 4 years of age were only obtained for a subset of the participants by design. Details on missing data are provided in Table S1 of Section 1 of the Supporting Information. Only participants with complete data (no missing data on any of the relevant variables, n=252) were included in the case study, as evaluation of challenges associated with the presence of missing data were not considered here. Characteristics of the infants included are outlined in Table LABEL:table:BIS_descriptive. Here, we have clearly identified a problem of high-dimensional confounding because we have a sample size of 252 and up to 87 potential confounders, with 78 of these being continuous confounders.
Figure 1: Directed Acyclic Graph (BIS DAG) for the Barwon Infant Study motivating example, and as used for data-generation in the simulation study.
Table 1: Characteristics of participants in the Barwon Infant Study (BIS) analytic sample used in the case study.
Inflammation^
Characteristic Low, N = 1891 High, N = 631 Overall, N = 2521
Pre-pregnancy BMI (kg/m2) 24.5 [21.7, 27.8] 23.5 [21.3, 28.4] 24.3 [21.7, 27.9]
Socio-Economic Indexes for Areas (SEIFA)
          Low 65 (34%) 19 (30%) 84 (33%)
          Med 63 (33%) 21 (33%) 84 (33%)
          High 61 (32%) 23 (37%) 84 (33%)
Mother smoking in pregnancy 22 (12%) 10 (16%) 32 (13%)
Gestational diabetes in pregnancy 8 (4.2%) 3 (4.8%) 11 (4.4%)
Pre-eclampsia in pregnancy 7 (3.7%) 2 (3.2%) 9 (3.6%)
Birthweight (grams) 3,531 (521) 3,442 (581) 3,508 (537)
Infant sex
          Female 82 (43%) 37 (59%) 119 (47%)
          Male 107 (57%) 26 (41%) 133 (53%)
Maternal age at birth (years) 32.1 (4.3) 32.8 (3.9) 32.3 (4.2)
Gestational age at birth
          32-36 completed weeks 9 (4.8%) 4 (6.3%) 13 (5.2%)
          37-42 completed weeks 180 (95%) 59 (94%) 239 (95%)
Mode of birth
          Caesarean 75 (40%) 20 (32%) 95 (38%)
          Vaginal 114 (60%) 43 (68%) 157 (62%)
Weight-for-length z-score at 12 months 0.72 (1.09) 0.75 (1.04) 0.72 (1.08)
Age at 12-month measures (months) 12.93 (0.80) 12.97 (0.75) 12.94 (0.79)
Breastfeeding duration (exclusive weeks*) 8 [1, 22] 4 [0, 24] 7 [0, 22]
Postnatal smoke exposure 29 (15%) 10 (16%) 39 (15%)
1Mean (SD), Median [IQR] or Frequency (%) as appropriate.
*Number of weeks that infant was exclusively breastfed (i.e. no supplementary feeding).
^GlycA at 1-year of age, dichotomised using the 75th percentile of the observed distribution as the threshold.
@@ 3.1 Causal estimand
We focus on the average causal effect (ACE), defined as $E[Y^{X=1}]-E[Y^{X=0}]$, interpreted as the difference in average potential outcome in the population of interest when (a) everyone is exposed versus when (b) everyone is unexposed. 36, 37 In the motivating example, this is the difference in average PWV that would be seen at 4 years of age if everyone in the target population were set to have high versus low GlycA at 1 year of age.
@@ 4 Doubly robust methods
Under the assumptions outlined in Section 3, the ACE can be estimated using DR methods. We focus on two DR methods, AIPW and TMLE. For both of these methods, models are fitted for the conditional expectation of the outcome $Y$ given the exposure $X$ and confounders $W$, $E[Y|X,W]$ (the outcome model) and for the propensity score, ${P}(X=1|W)$ (the exposure model). The outcome and exposure models are often referred to as nuisance models because they are not of intrinsic interest but instead are used within AIPW and TMLE to estimate the target parameter. 39
DR estimators are obtained by determining and then utilising an efficient influence function (EIF), which has a unique form that is determined by the target parameter of interest, in our case the ACE.22 DR estimators are constructed in a manner that allows them to attempt to correct for the bias (termed plug-in bias) that is induced when using g-computation with the data-adaptive approaches. Further details regarding derivation and interpretation of the EIF, as well as plug-in bias can be found elsewhere. 40
@@ 4.1 Augmented Inverse Probability Weighting (AIPW)
AIPW is often referred to as a one-step correction or one-step estimation approach 41 and is based on directly subtracting an estimate of the bias term from the g-computation estimator. Based on the fitted outcome model, predicted values, $\hat{E}_{1}(w)=\hat{E}[Y|X=1,W=w]$ and $\hat{E}_{0}(w)=\hat{E}[Y|X=0,W=w]$ are obtained for each record, for $X=1$ and $X=0$ respectively. The estimate of the ACE is then calculated as
@@
where $\hat{P}(w)=\hat{P}(X=1|W=w)$ is an estimate of the propensity score based on the fitted exposure model. The estimated variance of the AIPW estimator is obtained as follows: 42, 23